# Entanglement signatures of a percolating quantum system

### v1

@@ Abstract
Entanglement measures have emerged as one of the versatile probes to diagnose quantum phases and their transitions. Universal features in them expand their applicability to a range of systems, including those with quenched disorder. In this work, we show that when the underlying lattice has percolation disorder, free fermions at a finite density show interesting entanglement properties. These emerge from massively degenerate eigenspaces where the notion of standard bipartite entanglement entropy needs to be revisited. We show that appropriately defined entanglement measures can capture both geometrical aspects and electronic correlations of the percolated quantum system. Using a variety of toy models, we uncover the rich phenomenology of this physics. We further show that there are decipherable signatures of the classical percolation transition in these entanglement properties. Our work presents an interesting interplay between the geometrical properties of a lattice and quantum entanglement in a many-body quantum system.
@@
Introduction: Quantum entanglement [1, 2] and their measures [3, 4, 5] have established themselves as necessary tools to diagnose quantum phases and their transitions in a variety of systems, be it clean or disordered[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The universal features in quantities such as bipartite entanglement entropy [3] and entanglement spectrum [17] have also made it a defining quantity for studying unconventional quantum phases such as topological phases [18, 19, 17, 20, 21, 22, 23], particularly in presence of disorder [24, 25]. Among disordered systems, percolation problems [26, 27, 28, 29], where a lattice is probabilistically diluted either on the sites or bonds, are known to exhibit second-order phase transitions, namely geometrical phase transition, with universal critical exponents [30, 31, 32, 29, 33]. The critical point is known as the classical percolation threshold $p_{c}$, where the lattice suddenly gets geometrically disconnected [29]. The interplay of such geometric transitions has been long studied in both classical and quantum systems [34, 35, 36, 37, 38, 39] and recently in the context of topological phases [40, 41, 42]. However, the interplay of percolation disorder and entanglement properties has been little explored [43]. We visit entanglement measures in light of the percolation disorder in the simplest of fermionic quantum systems: free fermions hopping on a lattice. In particular, we address do entanglement measures show signatures of a geometrical phase transition? Do they still follow the usual area/volume law [44, 45] diagnostics?
@@
In this work, we investigate above questions to show that in percolating quantum systems, the conventional notions of bipartite entanglement properties do not hold. Since we will focus on the free fermionic systems, it is expected that entanglement entropy should be straightforwardly calculated using the Peschel prescription of correlator matrices [46, 47]. However, as we will show in percolation disorder, such calculations have quite a few subtleties. The non-intuitive aspect of the results, particularly in entanglement, arises from massive exact degeneracy due to lattice percolation. In particular, we show, that one needs to investigate three distinct quantities: typical entanglement ($S_{\text{typ}}$), annealed entanglement ($S_{\text{ann}}$) and quenched entanglement ($S_{\text{quen}}$) each of which captures distinctive signatures. Interestingly, they often depend on the number of exact zero-modes ($N_{0}$) in the system, which is in turn related to the geometrical aspects of the lattice. We finally show that even the classical percolation threshold has distinctive signatures, where entanglement scaling depends on the emergent fractal nature of the largest cluster.
@@
Measures of entanglement and free fermions: Given a wavefunction $|\psi\rangle$, or the density matrix ($\hat{\rho}=|\psi\rangle\langle\psi|$) defined for a system composed of two subsystems $A$ and $B$, the entanglement entropy of region $A$ with $B$ (or vice-versa) is given by $S_{A}=-\text{Tr}\big{(}{\hat{\rho}_{A}\log\hat{\rho}_{A}}\big{)}$, where $\hat{\rho}_{A}=\text{Tr}_{B}\hat{\rho}$. However, this definition assumes that the system is described by a unique wavefunction $|\psi\rangle$. But in presence of degeneracies, it is more appropriate to study typical entanglement
@@
where the averaging on the reduced density matrix is done before evaluating its entanglement content. Such a measure we call annealed entanglement entropy, where the annealed average renders $\langle\hat{\rho}_{A}\rangle$ to be of a mixed character. In general $S_{\text{typ}}<S_{\text{ann}}$, since the latter reflects the maximum information content possible in a given Hilbert space consistent with Page’s result [48]. For instance, in a decoupled two spin system while $S_{A}$ may seem to be zero, $S_{\text{typ}}\sim 0.33$, $S_{\text{ann}}=\log(2)$ (for details see Supplemental Material (SM) [54]).
In a many-body free fermionic system defined on $N$ sites such that $c^{\dagger}_{i}$ and $c_{i}$ are the fermionic creation and annihilation operators, the ground state is often a unique Slater determinant and the bipartite entanglement is evaluated using the Peschel formulation of correlator matrix ${\cal C}$ where its elements $c_{ij}=\langle c^{\dagger}_{i}c_{j}\rangle$[46, 47]. However, for a system with ground state degeneracies, the measure of typical and annealed entanglement becomes important. Thus while for any single Slater determinant, the entanglement entropy can be given by the corresponding correlator matrix $S({\cal C}_{A})$, when averaged over all choices of such random coefficients over the degenerate manifold, one gets $S_{\text{typ}}=\langle S({\cal C}_{A})\rangle$ and similarly $S_{\text{ann}}=S(\langle{\cal C}_{A}\rangle)$ (for an illustrative example of these quantities in a four-site toy model, see SM [54]). A simple way to break such exact degeneracies would be to add a small perturbative Anderson disorder $\eta\sim 10^{-12}$, which will choose a unique ground state. Then we define quenched entanglement, to capture the fermionic correlations in the system
@@
Figure 1: Zero dimensional model: (a) $S_{\text{quen}}$ (averaged over $N_{C}\equiv 10^{3}$ configurations) grows linearly with number of sites, $N$. (inset) Schematic of the model and partitions $A$ and $B$. (b) Behaviour of $S_{\text{typ}}$ and $S_{\text{ann}}$ with $N$. The number of zero-modes, $N_{0}=N-1$. Both are calculated using $N_{R}\equiv 10^{4}$ ensembles of random pure states.
@@
Zero-dimensional system: Consider a $N$ site Hamiltonian where the fermions can hop from a site to any other site with strength $t=-1$ (see Fig. 1(a) inset) such that $H=-\sum_{i,j}c^{\dagger}_{i}c_{j}$. The single particle spectrum has one eigenvalue with energy $-N$, while $N-1$ eigenvalues are exactly $zero$. A half-filled system here is thus ${}^{N-1}C_{\frac{N}{2}-1}$ fold degenerate. An infinitesimal $\eta$ disorder of strength $\eta=10^{-12}$ can be added in form of a term $\sum_{i}\epsilon_{i}c^{\dagger}_{i}c_{i}$ where $\epsilon\in[-\eta,\eta]$ to break this exact degeneracy. Given the long-range character of hopping, the entanglement content for equal bipartition of the system such that $N_{A}=N/2$ is still volume law where $S_{\text{quen}}=s_{v}N+s_{0}$ with a coefficient $s_{v}\sim 0.002$ as shown in Fig. 1(a). However, when such massive degeneracy is intact - that reflects in
$S_{\text{typ}}=S_{\text{quen}}+(\log 2-\frac{1}{2})(N_{0}-1),$ (4)
@@
where the second term arises from the effective geometrical component because of $N_{0}=N-1$ number of zero-modes of the system (see Fig. 1(d)). The factor of $(\log 2-\frac{1}{2})$ arises from the effective random pure states made out of $N_{0}$ zero-modes [55]. In general, when the number of occupied zero-modes is $fN_{0}$ ($10<f<$ 1), the geometrical volume law is empirically proportional to $f(1-f)$ representing the effective phase space volume (see SM [54]). Another alternate measure of the entanglement in this system is to average the correlator matrix first, which in general, corresponds to a mixed-density matrix. For such a system while $c_{ii}=\frac{1}{2}$ and $c_{ij}=\frac{1}{2N}$ [56] leading to $S_{\text{ann}}\sim\frac{N_{0}}{2}\log 2.$
@@
While it may appear that all entanglement measures are volume law and therefore similar, it is pertinent to emphasize that all three quantities represent three distinct entanglement measures having different physical content. While $S_{\text{quen}}$ captures the fermionic correlations, which are inherently volume law since the network is zero-dimensional, the character of $S_{\text{typ}}$ and $S_{\text{ann}}$ capture the massive degeneracy of the system - where $S_{\text{typ}}$ measures the average entanglement for any choice of a typical pure state, $S_{\text{ann}}$ measures the entanglement due to a mixed state arising from averaging the density matrix first. For instance, the entanglement measure of the full system has $S_{\text{ann}}\neq 0$ but $S_{\text{typ}}=0$. Having discussed the subtleties and the different measures of entanglement, we now discuss quantum percolation problems, where studying these various measures of entanglement become indispensable given spectral degeneracies.
@@
One-dimensional percolation: We first discuss percolation in one-dimensional lattice, where spinless fermions hop with the following tight-binding Hamiltonian
@@
where $L$ is the system size and $t=1$. The probability of having a bond is given by $p$ such that at $p=0$, the system contains a completely decoupled set of sites, while at $p=1$ it is a translationally invariant one-dimensional chain (see Fig. 2(a)). The percolation transition, where the connectivity of one end of the lattice gets disconnected from the other end, happens at $p=p_{c}=1$ [31, 57]. The fermion filling is kept fixed at $=1/2$. Given the Hamiltonian is real and has a sublattice symmetry, it belongs to the BDI symmetry class [58, 59], which is retained under the percolation protocol.
@@
The fermionic ground state describes a Fermi sea at $p=1$; however, at any $p<1$, given the fermions reside on disconnected clusters, the ground state should be interpreted as an Anderson insulator state. This is consistent with the effect of any uncorrelated disorder in one dimension[60, 61]. The entanglement content, therefore, is generically expected to be $\sim\log L$ at $p=1$ and $\mathcal{O}(1)$ in presence of disorder, as is known from Cardy-Calabrese result for critical states [10, 62, 12] and area law entanglement for short-range correlated states [44, 14]. At any $p$ given the presence of disconnected clusters, there are spectral degeneracies that can be split using an infinitesimal disorder $\eta$ to obtain $S_{\text{quen}}$ (see Fig. 2(b)). An analytical estimate can be obtained as follows. Given $P_{s}$ is the probability of having a $s$ cluster, in general, its maximal entanglement content in equal bi-partition is ($\frac{c}{6}\log(\frac{s}{\pi})+c_{0}$) [10] where $c$ is the central charge ($c=1$) of the one-dimensional bosonic CFT and $c_{0}$ is an area law piece. Thus for a thermodynamic system,
$\tilde{S}_{\text{quen}}=\sum_{s=2}^{\infty}P_{s}\Big{(}\frac{c}{6}\log(\frac{s}{\pi})+c_{0}\Big{)}$ (6)
@@
where $\tilde{S}_{\text{quen}}$ represents a configuration averaged value over $S_{\text{quen}}$. Here $P_{s}=sn_{s}=s(1-p)^{2}p^{s-1}$ which, given any $p$, is the probability that an arbitrary site of the diluted chain belongs to a cluster of $s$ number of sites [29]. This has a linear rise at small $p$ and a divergence near $p=1$. This analytic behaviour, along with our numerical results, is shown in Fig. 2(b)(here, $c_{0}=0.409\pm 0.002$). Broadly the behaviour indeed remains area law (with a $p$ dependent coefficient) except at $p=1$ where the logarithmic $L$ dependence is restored. Interestingly the average cluster size $\langle M\rangle=\sum_{s}sP_{s}$, diverges near $p\rightarrow p_{c}$ as $(p_{c}-p)^{-\gamma}$ with $\gamma=1$. Thus $\tilde{S}_{\text{quen}}\sim\frac{c}{6}\log\langle M\rangle\sim\frac{c}{6}\log\xi^{\frac{\gamma}{\nu}}$, where the geometric correlation length $\xi$ also diverges as $(p_{c}-p)^{-\nu}$ with $\nu=1$. Since at $p=1$, $\xi\sim L$, one expects a scaling where, $e^{6\tilde{S}_{\text{quen}}}L^{-\frac{\gamma}{\nu}}\sim 1$ which is shown in Fig. 2(c). Thus, the entanglement measure captures the percolation exponents near the geometrical phase transition here at $p_{c}=1$. However, these results, as discussed before, required us to put an infinitesimal degeneracy splitting disorder $\eta$, which also breaks the symmetry of the full Hamiltonian. In absence of any such disorder, the massive degenerate manifold of zero-modes leads to geometrical components of entanglement entropy, as we discuss next.
@@
for a $L$ sized chain at percolation probability $p$. This analytical behaviour and the disorder averaged numerical estimate of zero-modes number $\tilde{N_{0}}$ are shown in Fig. 2(e). Now, a half-filled state in such a system again leads to highly degenerate many-body eigenspace. A uniform Haar measure here leads to
$\tilde{S}_{\text{typ}}=\tilde{S}_{\text{quen}}+(\log 2-\frac{1}{2})\tilde{N}_{0},$ (8)
while $\tilde{S}_{\text{ann}}\sim\frac{\tilde{N}_{0}}{2}\log 2$. All the behaviours match the numerical results as shown in Fig. 2(d). Thus, geometrical disorder, as in one-dimensional percolation, provides distinctive signatures in various entanglement measures both from intra-cluster fermionic correlations and from geometric components of the lattice itself.
@@
Two-dimensional percolation: In two dimensional percolation the system has a finite $p_{c}$; for instance in square lattice bond percolation it is known $p_{c}=\frac{1}{2}$ [64, 65]. At $p=p_{c}$ a spanning cluster develops with critical exponent $\gamma=43/18$ [29]. We again pose the question of different entanglement measures for this system. As is known that for a two-dimensional free fermionic system - any infinitesimal disorder localizes all the wavefunctions [60, 61], thus in terms of electronic properties, we expect the system to be localized for all values of $p$ [66, 67, 68, 69] even though there have been studies finding numerical evidence otherwise [70, 71, 72, 73, 74, 75, 76]. Given any $p$ the mean number of clusters containing $k$ bonds, $n_{k}$ is given by $n_{k}=\sum_{t}g(k,t)p^{k}(1-p)^{t}$ where $t$ is the perimeter of the cluster and $g(k,t)$ is the geometrical factor associated with the number of lattice animals given ($k,t$) [77]. Since any square lattice is a bipartite graph with a symmetric spectrum, the $E_{F}$ remains pinned to zero even under percolation. We find that $S_{\text{quen}}$ follows an area law behaviuor (see Fig. 3(a)) i.e. $\propto L$ for $p<1$. While at $p=1$, the complete square lattice is restored, leading to a finite Fermi sea, the entanglement is proportional to $L\log L$ given by the Widom conjecture [12, 78]. At $p<1$, however, such a momentum space description is no longer applicable. At small $p$, one can enumerate the lattice animals exactly and count their entanglement contribution, as shown by the analytical curve in Fig. 3(a) (for details, see SM [54]). This is in contrast to $S_{\text{typ}}$ and $S_{\text{ann}}$, which again depend on the extensive number of zero-modes present in the system (see Fig. 3(b)). The density of zero-modes can again be estimated analytically from lattice animals, given by
@@
To distil any signature of percolation transition at $p=p_{c}$, we calculate the configuration averaged (quenched) bipartite entanglement entropy of the largest cluster ($\equiv S^{lc}_{\text{quen}}$) as a function of $p$ and show this in Fig. 4(a). Interestingly, while for $p<p_{c}$ $S^{lc}_{\text{quen}}\propto L^{0}$, for $p>p_{c}$ $S^{lc}_{\text{quen}}\propto L$ where $L$ is the linear dimension of the lattice. Given it is known that the largest cluster follows a scaling $L^{d_{f}}$ [29] where $d_{f}$ is the fractal dimension of the system, for a typical area law behaviour, we expect $S^{lc}_{\text{quen}}\propto L^{d_{f}/2}$ near $p_{c}$. A scaling collapse using this form shows a crossing (see Fig. 4(b)) at $p=p_{c}=0.5$ with $d_{f}=91/48$ as known for two-dimensional percolation transition. Since this physics should be independent of the microscopic lattice, we apply the same analysis to a tight-binding triangular lattice (Fig. 4(c)). Again, the same collapse shows a crossing (Fig. 4(d)) near $p=p_{c}=2\sin(\frac{\pi}{18})$ [64] illustrating that the entanglement scaling indeed follows the universal features of geometrical phase transitions even though the exact value of $p_{c}$ is itself not universal.

### v2

@@ Abstract
Entanglement measures have emerged as one of the versatile probes to diagnose quantum phases and their transitions. Universal features in them expand their applicability to a range of systems, including those with quenched disorder. In this work, we show that when the underlying lattice has percolation disorder, free fermions at a finite density show interesting entanglement properties due to massively degenerate ground states. We define and calculate appropriate entanglement measures such as typical, annealed, and quenched entanglement entropy in both one and two dimensions, showing they can capture both geometrical aspects and electronic correlations of the percolated quantum system. In particular, while typical and annealed entanglement show volume law character directly dependent on the number of zero modes in the system, quenched entanglement is generally area law albeit showing characteristic signatures of the classical percolation transition. Our work presents an exotic interplay between the geometrical properties of a lattice and quantum entanglement in a many-body quantum system.
@@
Introduction: Quantum entanglement [1, 2] and their measures [3, 4, 5] have established themselves as necessary tools to diagnose quantum phases and their transitions [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Moreover, the generalizations and extensions of bipartite entanglement entropy (EE) [3] including topological EE [17, 18], entanglement negativity [19, 20, 21, 22, 23, 24], witnesses [25, 26, 27, 28] and corner contributions [29, 30, 31, 32, 33] provide an encompassing framework to study quantum phases in and out of equilibrium. Their universal features at times make them indispensable to define unconventional quantum phases [34, 35, 36, 37, 38], particularly in the presence of disorder [39, 40]. Among disordered systems, percolation problems [41, 42, 43], where a lattice is probabilistically diluted either on the sites or bonds, are known to exhibit second-order phase transitions, namely geometrical phase transition, with universal critical exponents [44, 45, 46, 43, 47]. For instance, in the case of bond percolation if $p$ is the probability of having a bond in the system such that $p=1$ is a translationally invariant lattice then there exists a critical value of $p$ known as the classical percolation threshold $p_{c}$, immediately below which the lattice gets geometrically disconnected [43]. Such geometrical phase transitions and their critical phenomena have been long studied in both classical and quantum systems [48, 49, 50, 51, 52, 53] and recently in the context of topological phases [54, 55, 56]. However, the interplay of percolation disorder and entanglement properties has been little explored [57]. We visit entanglement measures in the light of percolation disorder in the simplest of fermionic quantum systems: free fermions hopping on a lattice. In particular, we address do entanglement measures show signatures of a geometrical phase transition? Do they still follow the usual area/volume law [58, 59] diagnostics?
@@
In this work, we investigate the above questions to show that in percolating quantum systems, the conventional measures of bipartite entanglement entropy [60, 61] have quite a few subtleties. The non-intuitive aspect of the results arises from massive exact degeneracy due to lattice percolation. In particular, we show that one needs to investigate different quantities: typical ($S_{\text{typ}}$), annealed ($S_{\text{ann}}$) and quenched EE ($S_{\text{quen}}$) each of which captures distinctive signatures. Interestingly, they often depend on the number of exact zero-modes ($N_{0}$) in the system, which is, in turn, related to the geometrical aspects of the lattice. We finally show that even the classical percolation threshold has footprints in quantum entanglement, where entanglement scaling depends on the emergent fractal nature of the largest cluster.
@@
Measures of entanglement and free fermions: Given a wavefunction $|\psi\rangle$, or the density matrix ($\hat{\rho}=|\psi\rangle\langle\psi|$) of a system composed of two subsystems $A$ and $B$, the EE of region $A$ with $B$ (or vice-versa) is given by $S_{A}=-\text{Tr}\big{(}{\hat{\rho}_{A}\ln\hat{\rho}_{A}}\big{)}$, where $\hat{\rho}_{A}=\text{Tr}_{B}\hat{\rho}$ is the reduced density matrix of the subsystem $A$. However, this definition assumes that the system is described by a unique wavefunction $|\psi\rangle$. But in the presence of degeneracies, it is more appropriate to study typical entanglement
@@
where the averaging is done on the density matrix before evaluating its entanglement content. Such a measure we call annealed EE. In general $S_{\text{typ}}<S_{\text{ann}}$, since the latter reflects the maximum information content possible in a given Hilbert space consistent with Page’s result [62]. For instance, in a decoupled two spin system while $S_{A}$ may seem to be zero, $S_{\text{typ}}\sim 0.33$, $S_{\text{ann}}=\ln 2$ (see Supplemental Material (SM) [68]).
In a many-body free fermionic system defined on $N$ sites such that $c^{\dagger}_{i}$ and $c_{i}$ are the fermionic creation and annihilation operators, the ground state is often a unique Slater determinant and the bipartite entanglement is evaluated using the Peschel formulation of correlator matrix ${\cal C}$ where its elements $c_{ij}=\langle c^{\dagger}_{i}c_{j}\rangle$ [60, 61]. However, the measure of typical and annealed entanglement becomes important for a system with ground state degeneracy. Thus, while for any single Slater determinant, the EE can be given by the corresponding correlator matrix $S({\cal C}_{A})$, when averaged over all choices of random wavefunction coefficients over the degenerate manifold, one gets $S_{\text{typ}}=\langle S({\cal C}_{A})\rangle$ and similarly $S_{\text{ann}}=S(\langle{\cal C}_{A}\rangle)$ (for an illustrative example of these quantities in a four-site toy model, see SM [68]). A simple way to break such exact degeneracies would be to add a small perturbative Anderson disorder $\eta\sim 10^{-12}$, which will choose a unique ground state. Then we define quenched entanglement, to capture the fermionic correlations in the system
@@
Figure 1: Zero dimensional model: (a) $S_{\text{quen}}$ (averaged over $N_{C}\equiv 10^{3}$ configurations) grows linearly with number of sites, $N$. (inset) Schematic of the model and subsystems $A$ and $B$. (b) Behaviour of $S_{\text{typ}}$ and $S_{\text{ann}}$ with $N$. The number of zero-modes, $N_{0}=N-1$. Both are calculated using $N_{R}\equiv 10^{4}$ ensembles of random pure states.
@@
Zero-dimensional system: Consider a $N$ site Hamiltonian where the fermions can hop from a site to any other site with strength $t=-1$ (see Fig. 1(a) inset) such that $H=-\sum_{i,j}c^{\dagger}_{i}c_{j}$. The single particle spectrum has one eigenvalue with energy $-N$, while $N-1$ eigenvalues are exactly $zero$. A half-filled system here is thus (${N-1}\choose{\frac{N}{2}-1}$ fold degenerate. An infinitesimal $\eta$ disorder of strength $\eta=10^{-12}$ can be added in form of a term $\sum_{i}\epsilon_{i}c^{\dagger}_{i}c_{i}$ where $\epsilon\in[-\eta,\eta]$ to break this exact degeneracy. Given the long-range character of hopping, the entanglement content for equal bipartition of the system such that $N_{A}=N/2$ is still volume law where $S_{\text{quen}}=s_{v}N+s_{0}$ with a coefficient $s_{v}\sim 0.004$ as shown in Fig. 1(a). However, when such massive degeneracy is intact, reflects in
$S_{\text{typ}}=S_{\text{quen}}+(\ln 2-\frac{1}{2})(N_{0}-1),$ (4)
@@
where the second term arises from the effective geometrical component because of $N_{0}=N-1$ number of zero-modes of the system (see Fig. 1(b)). The factor of $(\ln 2-\frac{1}{2})$ arises from the effective random pure states made out of $N_{0}$ zero-modes [69]. In general, when the number of occupied zero-modes is $fN_{0}$ ($10<f<$ 1), the geometrical volume law is empirically proportional to $f(1-f)$ representing the effective phase space volume (see SM [68]). Another alternate measure of the entanglement in this system is to average the correlator matrix first. For such a system at half-filling, while $c_{ii}=\frac{1}{2}$ and $c_{ij}=\frac{1}{2N}$ [70] leading to $S_{\text{ann}}\sim\frac{N_{0}}{2}\ln 2$.
@@
While it may appear that all three entanglement measures are volume law and therefore similar, it is pertinent to emphasize that they all have different physical content. While $S_{\text{quen}}$ captures the fermionic correlations, which are inherently volume law since the network is zero-dimensional, the character of $S_{\text{typ}}$ and $S_{\text{ann}}$ capture the massive degeneracy of the system - where $S_{\text{typ}}$ measures the average bipartite EE for any choice of a typical pure state, $S_{\text{ann}}$ measures the maximal subsystem entanglement when the complete system itself becomes mixed due to averaging of the correlator matrix. For instance, the entanglement measure of the full system has $S_{\text{ann}}\neq 0$ but $S_{\text{typ}}=0$. In general, when a complete system of interest takes a mixed character various other entanglement measures have been found to isolate quantum correlations between its partitions such as mutual information [71, 72, 73], entanglement negativity [19, 20, 21, 22, 23, 24] and witnesses [25, 26, 27, 28]. Having discussed the subtleties and the different measures of entanglement, we now discuss quantum percolation problems, where studying these various measures of entanglement becomes indispensable given spectral degeneracies.
@@
One-dimensional percolation: We first discuss percolation in a one-dimensional lattice, where spinless fermions hop with the following tight-binding Hamiltonian
@@
where $L$ is the system size and $t=1$. The probability of having a bond is given by $p$ such that at $p=0$, the system contains a completely decoupled set of sites, while at $p=1$ it is a translationally invariant fermionic chain (see Fig. 2(a)). The percolation transition, where one end of the lattice gets connected to the other end, happens at $p=p_{c}=1$ [45, 74]. The fermion filling is kept fixed at $=1/2$. Given the Hamiltonian is real and has a sublattice symmetry, it belongs to the BDI symmetry class [75, 76], which is retained under the percolation protocol.
@@
The fermionic ground state describes a Fermi sea at $p=1$; however, at any $p<1$, given the fermions reside on disconnected clusters, the ground state should be interpreted as an Anderson insulator state. This is consistent with the effect of any uncorrelated disorder in one dimension[77, 78]. The entanglement content, therefore, is generically expected to be $\sim\ln L$ at $p=1$ and $\mathcal{O}(1)$ in presence of disorder, as is known from Cardy-Calabrese result for critical states [10, 79, 12] and area law entanglement for short-range correlated states [58, 14]. At any $p$ given the presence of disconnected clusters, there are spectral degeneracies that can be split using an infinitesimal disorder $\eta$ to obtain $S_{\text{quen}}$ (see Fig. 2(b)). An analytical estimate can be obtained as follows. Given $P_{s}$ is the probability of having a $s$ sized cluster, in general, its maximal entanglement content in equal bipartition is ($\frac{c}{6}\ln(\frac{s}{\pi})+c_{0}$) [10] where $c$ is the central charge ($c=1$) of the one-dimensional bosonic CFT and $c_{0}$ is an area law piece. Thus for a thermodynamic system,
$\tilde{S}_{\text{quen}}=\sum_{s=2}^{\infty}P_{s}\Big{(}\frac{c}{6}\ln(\frac{s}{\pi})+c_{0}\Big{)}$ (6)
@@
where $\tilde{S}_{\text{quen}}$ represents a configuration averaged value over $S_{\text{quen}}$. Here $P_{s}=sn_{s}=s(1-p)^{2}p^{s-1}$ which, given any $p$, is the probability that an arbitrary site of the diluted chain belongs to a cluster of $s$ number of sites [43]. $n_{s}$ is the mean number of clusters of size $s$. This has a linear rise at small $p$ and a divergence near $p=1$. This analytic behavior, along with our numerical results, is shown in Fig. 2(b)(here, $c_{0}=0.409\pm 0.002$). Broadly, the behavior indeed remains area law (with a $p$ dependent coefficient) except at $p=1$ where the logarithmic $L$ dependence is restored. Interestingly, the average cluster size $\langle M\rangle=\sum_{s}sP_{s}$, diverges near $p\rightarrow p_{c}$ as $(p_{c}-p)^{-\gamma}$ with $\gamma=1$ [43]. Thus $\tilde{S}_{\text{quen}}\sim\frac{c}{6}\ln\langle M\rangle\sim\frac{c}{6}\ln\xi^{\frac{\gamma}{\nu}}$, where the geometric correlation length $\xi$ also diverges as $(p_{c}-p)^{-\nu}$ with $\nu=1$ [43]. Since at $p=1$, $\xi\sim L$, one expects a scaling where, $e^{6\tilde{S}_{\text{quen}}}L^{-\frac{\gamma}{\nu}}\sim 1$ which is shown in Fig. 2(c). Thus, the entanglement measure captures the percolation exponents near the geometrical phase transition here at $p_{c}=1$. However, as discussed before, these results required us to put an infinitesimal degeneracy splitting disorder $\eta$, which also breaks the symmetry of the full Hamiltonian. Without any such disorder, the massive degenerate manifold of zero-modes leads to geometrical components of EE, as we discuss next.
@@
for a $L$ sized chain at percolation probability $p$. This analytical behavior and the disorder averaged numerical estimate of zero-modes number $\tilde{N_{0}}$ are shown in Fig. 2(e). A half-filled state in such a system again leads to highly degenerate many-body eigenspace. A uniform Haar measure here leads to
$\tilde{S}_{\text{typ}}=\tilde{S}_{\text{quen}}+(\ln 2-\frac{1}{2})\tilde{N}_{0},$ (8)
while $\tilde{S}_{\text{ann}}\sim\frac{\tilde{N}_{0}}{2}\ln 2$. All the behaviors match the numerical results as shown in Fig. 2(d). Interestingly, the mutual information between the two subsystems removes this large volume law contribution and shows a rise similar to $\tilde{S}_{\text{quen}}$, only near $p=1$ when the lattice gets connected (see SM [68]). Thus, geometrical disorder, as in one-dimensional percolation, provides distinctive signatures in various entanglement measures both from intra-cluster fermionic correlations and from geometric components of the lattice itself.
@@
Two-dimensional percolation: In two dimensional percolation the system has a finite $p_{c}$; for instance in square lattice bond percolation it is known $p_{c}=\frac{1}{2}$ [81, 82, 83]. At $p=p_{c}$ a spanning cluster develops with critical exponent $\gamma=43/18$ [43]. We again pose the question of different entanglement measures for this system. As is known that for a two-dimensional free fermionic system - any infinitesimal disorder localizes all the wavefunctions [77, 78], thus in terms of electronic properties, we expect the system to be localized for all values of $p$ [84, 85, 86, 87] even though there have been studies finding numerical evidence otherwise [88, 89, 90, 91, 92, 93, 94]. Given any $p$ the mean number of clusters containing $k$ bonds, $n_{k}$ is given by $n_{k}=\sum_{t}g(k,t)p^{k}(1-p)^{t}$ where $t$ is the perimeter of the cluster and $g(k,t)$ is the geometrical factor associated with the number of lattice animals given ($k,t$) [95]. Since any square lattice is a bipartite graph with a symmetric spectrum, the $E_{F}$ remains pinned to zero at half-filling even under percolation. We find that $S_{\text{quen}}$ follows an area law behavior (see Fig. 3(a)) i.e. $\propto L$ for $p<1$. While at $p=1$, the complete square lattice is restored, leading to a finite Fermi sea, the entanglement is $\sim L\ln L$ given by the Widom conjecture [12, 96]. At $p<1$, however, such a momentum space description is no longer applicable. At small $p$, one can enumerate the lattice animals exactly and count their entanglement contribution, as shown by the analytical curve in Fig. 3(a) (for details, see SM [68]). This is in contrast to $S_{\text{typ}}$ and $S_{\text{ann}}$, which again depend on the extensive number of zero-modes present in the system (see Fig. 3(b)). The density of zero-modes can be estimated analytically from lattice animals, given by
@@
To distill any signature of percolation transition at $p=p_{c}$, we calculate the configuration averaged quenched bipartite EE of the largest cluster ($\equiv S^{lc}_{\text{quen}}$) as a function of $p$ and show this in Fig. 4(a). Interestingly, while for $p<p_{c}$ $S^{lc}_{\text{quen}}\propto L^{0}$, for $p>p_{c}$ $S^{lc}_{\text{quen}}\propto L$ where $L$ is the linear dimension of the lattice. Given it is known that the largest cluster follows a scaling $L^{d_{f}}$ [43] where $d_{f}$ is the fractal dimension of the system, for a typical area law behavior, we expect $S^{lc}_{\text{quen}}\propto L^{d_{f}/2}$ near $p_{c}$. A scaling collapse using this form shows a crossing (see Fig. 4(b)) at $p=p_{c}=0.5$ with, $d_{f}=91/48$ as known for two-dimensional percolation transition. Since this physics should be independent of the microscopic lattice, we apply the same analysis to a tight-binding triangular lattice (Fig. 4(c)). Again, the same collapse shows a crossing (Fig. 4(d)) near $p=p_{c}=2\sin(\frac{\pi}{18})$ [82] illustrating that the entanglement scaling indeed follows the universal features of geometrical phase transitions even though the exact value of $p_{c}$ is itself not universal.